# A Compartmental Model for Epidemiology with Human Behavior and Stochastic Effects

### v2

@@ 4. Asymptotic behavior of the stochastic system near disease free states
(4.47) $C=\left(\tfrac{2\mu s^{2}+4\delta}{\nu}+\tfrac{6\beta+3\mu}{2\mu\varepsilon\delta}+\tfrac{8(\gamma^{2}+(\gamma+2\delta)^{2}+\beta^{2}(\delta/\nu))}{\varepsilon\delta}\right)\left(\tfrac{b}{\delta}\right)^{4}$
which has not been optimized in any way, but still demonstrates how the constant should behave under different parameter regimes. The only non-explicit part of the constant is $\varepsilon$ which is proportional the absolute value of the maximum eigenvalue of $F_{s,s^{*}}-V_{s,s^{*}}$. The eigenvalue itself is guaranteed to be negative when $\mathscr{R}(s,s^{*})<1$, but empirically, it will be closer to zero when $\mathscr{R}(s,s^{*})$ is closer to 1 and farther from zero when $\mathscr{R}(s,s^{*})$ gets much less than 1. Thus $C$ shrinks as $\mathscr{R}(s,s^{*})$ gets smaller and blows up when $\mathscr{R}(s,s^{*})\to 1$. Other things to notice is that $C\to\infty$ if $\nu\to 0,\mu\to 0$ or $\delta\to 0$ which makes intuitive sense. If $\nu\to 0$ there is no recovery from noncompliance so everyone becomes noncompliant and we can no longer guarantee that $(S-s)$ remains small; if $\mu\to 0$, then no one is becoming noncompliant so this population dies out and we cannot guarantee that $(S^{*}-s^{*})$ remains small; and as $\delta\to 0$, the upper bound on the population gets very large and thus the variance in the stochastic system also increases.
@@ 5. Simulations and discussion
We present some simulations to demonstrate various aspects of our model, and empirically verify the theorems. All simulations were performed in MATLAB. System (2.2) was discretized using a simple explicit Euler scheme, and (3.1) was discretized using Milstein’s scheme [19, Ch. 10].
@@
In our first simulation, we take $\beta=1$, $\xi=0$, $\mu=\nu=0.2$, and $\sigma_{\beta}=\sigma_{\mu}=0.5$. Plugging in these values, we find that $21.625=\tfrac{b}{\delta}+\frac{\sigma_{\mu}^{2}}{2\mu}\Big{(}\tfrac{b}{\delta}\Big{)}^{2}<\frac{\nu+\delta}{\mu}=$ 2 and $\mathscr{R}_{0}(\tfrac{b}{\delta},0)<\mathscr{R}^{\sigma}_{0}(\tfrac{b}{\delta})\approx 0.860.$ So the hypotheses of both theorem 2.5(i) and theorem 4.2 are met and the complaint disease free equilibrium $s=\tfrac{b}{\delta},s^{*}=0$ is stable. Indeed, this is born out in figure 1, where $S^{*},I,I^{*}$ all seemingly exhibit immediate exponential decay and $S$ simply increases to $s=\tfrac{b}{\delta}=1$.
@@
For our second simulation, we take $\beta=1,\xi=0$, $\mu=\nu=0.2$, and $\sigma_{\beta}=\sigma_{\mu}=2$. Here, the deterministic part is the same as in the first simulation, but we have substantially increased the randomness. In this case, $21=\tfrac{b}{\delta}<\tfrac{\nu+\delta}{\mu}=$ 2 and $\mathscr{R}_{0}(\tfrac{b}{\delta},0)\approx 0.803$ so the conditions of theorem 2.5(i) are met, and stability is maintained in the deterministic case, but $211=\tfrac{b}{\delta}+\frac{\sigma_{\mu}^{2}}{2\mu}\Big{(}\tfrac{b}{\delta}\Big{)}^{2}>\frac{\nu+\delta}{\mu}=$ 2 and $\mathscr{R}_{0}^{\sigma}(\tfrac{b}{\delta},0)\approx 1.708$ so neither of the conditions for theorem 4.2 are met and we cannot guarantee stability in the stochastic case. In this case, the stochastic simulations often exhibit nontrivial outbreaks of the disease as seen in figure 2, where there is an initial outbreak of the disease among the noncompliant population, and later a significant outbreak among the compliant population that correlates with a random spike in noncompliance. Interestingly, in all realizations we observed, by time $T=50$, the stochastic simulation seems to match the deterministic one, which gives empirical evidence that the sufficient conditions for stability in theorem 4.2 may not be necessary (though it is not clear how they could be relaxed in our proof).
@@
For our third simulation, we consider $\beta=0.6,\xi=1$, $\mu=0.1$, $\nu=0$, and $\sigma_{\beta}=\sigma_{\mu}=0.4$. In this case, $\mathscr{R}_{0}^{\sigma}(0,\tfrac{b}{\delta})\approx 0.971$ and $0.20.16=\Big{(}\tfrac{\beta}{\delta}\Big{)}^{2}\sigma_{\mu}^{2}<\delta=$ 0.2, so the conditions of theorem 4.4 are met (as well as those of theorem 2.5(iii) in the worst case scenario of $\xi=1$ and $\nu=0$), so the disease free equilibrium $s=0,s^{*}=\tfrac{b}{\delta}$ is stable for both the deterministic and stochastic systems. This is seen in figure 3. If however, we increase $\sigma_{\mu}=2$ (leaving all other parameters as specified), we still have $\mathscr{R}_{0}^{\sigma}(0,\tfrac{b}{\delta})\approx 0.971$, but now $0.24=\Big{(}\tfrac{\beta}{\delta}\Big{)}^{2}\sigma_{\mu}^{2}>\delta=$ 0.2. Thus according to theorem 4.4, the infections should still exhibit exponential decay, but we can no longer guarantee stability for the DFE. This is seen in figure 4 where the infections do indeed die out, but there is some initial fighting between $S$ and $S^{*}$. As mentioned in remark 4.5, this would be unlikely to cause any concern for the policy-maker, since the important consideration is whether or not infections persist, not the particular concentration of compliance and noncompliance once infections have died out, and any additional compliant population is actually a boon since they will reduce the effective reproductive ratio.
@@
In our final simulation, we vary several of the parameters (including some that were fixed in figures 1-4). Specifically, we set the birth rate to $b=0.5$, the death rate to $\delta=1$, the infectivity rate of the disease to $\beta=0.5$, the recovery rate for the disease to $\gamma=0.25$, the infectivity rate for noncompliance to $\mu=6,$ the recovery rate from noncompliance to $\nu=1$, and the uncertainty in the infectivity rates to $\sigma_{\beta}=\sigma_{\mu}=0.075$. We take $\xi=0$ so that newly introduced members of the population are uniformly compliant. The reduction in infectivity is held at $\alpha=0.25$. With all these decisions, we consider the disease free state $s=\frac{\nu+\delta}{\mu}=\frac{1}{3}$ and $s^{*}=\tfrac{b}{\delta}-\frac{\nu+\delta}{\mu}=\frac{1}{6}$. In this case $\mathscr{R}_{0}(s,s^{*})\approx 0.144$ so we are well within the local asymptotic stability regime provided by theorem 2.5(ii). Thus the solution of the deterministic system should tend toward the disease free equilibrium. While the stochastic system will not be at equilibrium at this pair $(s,s^{*})$, we do satisfy the hypothesis of theorem 4.7, and taking $C$ from (4.47), we have $C(\sigma_{\beta}^{2}+\sigma_{\mu}^{2})\approx 0.0197$. So while the solution of the stochastic system will not be steady, it should stay within these error bars of the deterministic solution. This is seen to occur in figure 5 where we plot the deterministic solution along with three realizations of the stochastic system for times $t\geq 10$. We see that the stochastic realizations stay well within the guaranteed error bars, which are displayed in gray.
@@ 6. Concluding remarks
In this manuscript we analyze an ODE model for epidemiology incorporating human behavioral effects. We assume there are governmental protocols enacted to stunt the spread of the disease, but that there is a subpopulation which will not comply with these protocols, and that this noncompliance spreads via mass action parallel to the spread of the disease. While our model is inspired by prior work, we give a more complete analysis of the stability of disease free equilibria. We then extend the model by adding stochastic perturbations, specifically assuming that there is uncertainty in the disease transmission rate and in the transmission rate of noncompliant behavior. We prove almost sure long time existence and positivity for the stochastic model and establish stability conditions for two disease free equilibria by constructing suitable stochastic Lyapunov functions. Finally, we consider points which are disease free equilibria of the deterministic model but not the stochastic model, and we quantify how much the stochastic model may stray from these points in the parameter regimes where they are stable for the deterministic model. We demonstrate our results with simulations.
One avenue of future work would be the identification of and stability conditions for endemic equilibrium points (where the disease does not die out) for the deterministic system and quantification of the behavior of stochastic system near these points. Another would be the incorporation of spatial effects resulting in a system of stochastic partial differential equation. There are some deterministic results in this direction in [7, 29] but adding stochasticity would lead to nontrivial complications in the analysis. Additionally, the inclusion of nonlocal effects wherein the disease and/or noncompliance can spread nonlocally could lead to interesting models requiring significantly modified analysis. Nonlocal effects may be especially apt in modeling spread of behaviors, where social media, news coverage or a variety of other sources may be seen to influence behavior without direct contact. One last path forward would be to consider the governmental protocols as control variables and design a cost functional for the policy-maker so that this becomes a stochastic control problem.
@@ Acknowledgments
WW was partially supported by a Junior Faculty Fellowship from the University of Oklahoma and a Simons Foundation grant (No. 0007730).

### v4

@@ 4. Asymptotic behavior of the stochastic system near disease free states
(4.49) $C=\left(\tfrac{2\mu s^{2}+4\delta}{\nu}+48\|Q\|_{2}\tfrac{2\gamma^{2}+2(\gamma+2\delta)^{2}+\beta^{2}(2\delta/\nu)}{\delta}\bigg(1+\tfrac{4\|Q\|_{2}(b/\delta)^{2}(\beta+\mu)}{3\mu}\bigg)\right)\left(\tfrac{b}{\delta}\right)^{4}$
which has not been optimized in any way, but still demonstrates how the constant should behave under different parameter regimes. The only non-explicit part of the constant is $\|Q\|_{2}$ where $Q$ is the solution of the Sylvester equation $M^{t}Q+QM=-I$ where $M\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=F_{s,s^{*}}-V_{s,s^{*}}$. This matrix $Q$ is quaranteed to exist if and only if all eigenvalues of $F_{s,s^{*}}-V_{s,s^{*}}$ have negative real part which holds if and only if $\mathscr{R}_{0}(s,s^{*})<1$. Thus we expect that this terms causes $C$ to blow up when $\mathscr{R}(s,s^{*})\to 1$. Other things to notice is that $C\to\infty$ if $\nu\to 0,\mu\to 0$ or $\delta\to 0$ which makes intuitive sense. If $\nu\to 0$ there is no recovery from noncompliance so everyone becomes noncompliant and we can no longer guarantee that (Ss)(S-s) remains small; if $\mu\to 0$, then no one is becoming noncompliant so this population dies out and we cannot guarantee that $(S^{*}-s^{*})$ remains small; and as $\delta\to 0$, the upper bound on the population gets very large and thus the variance in the stochastic system also increases.
@@ 5. Simulations and discussion
We present some simulations to demonstrate various aspects of our model, and empirically verify the theorems. All simulations were performed in MATLAB. System (2.2) was discretized using a simple explicit Euler scheme, and (3.1) was discretized using Milstein’s scheme [30, Ch. 10].
@@
In our first simulation, we take $\beta=1$, $\xi=0$, $\mu=\nu=0.2$, and $\sigma_{\beta}=\sigma_{\mu}=0.5$. Plugging in these values, we find that $1.625=\tfrac{b}{\delta}+\frac{\sigma_{\mu}^{2}}{2\mu}\Big(\tfrac{b}{\delta}\Big)^{2}<\frac{\nu+\delta}{\mu}=2$ and $\mathscr{R}_{0}(\tfrac{b}{\delta},0)<\mathscr{R}^{\sigma}_{0}(\tfrac{b}{\delta})\approx 0.860.$ So the hypotheses of both theorem 2.5(i) and theorem 4.2 are met and the complaint disease free equilibrium $s=\tfrac{b}{\delta},s^{*}=0$ is stable. Indeed, this is born out in figure 3, where $S^{*},I,I^{*}$ all seemingly exhibit immediate exponential decay and $S$ simply increases to $s=\tfrac{b}{\delta}=1$.
@@
For our second simulation, we take $\beta=1,\xi=0$, $\mu=\nu=0.2$, and $\sigma_{\beta}=\sigma_{\mu}=2$. Here, the deterministic part is the same as in the first simulation, but we have substantially increased the randomness. In this case, $1=\tfrac{b}{\delta}<\tfrac{\nu+\delta}{\mu}=2$ and $\mathscr{R}_{0}(\tfrac{b}{\delta},0)\approx 0.803$ so the conditions of theorem 2.5(i) are met, and stability is maintained in the deterministic case, but $11=\tfrac{b}{\delta}+\frac{\sigma_{\mu}^{2}}{2\mu}\Big(\tfrac{b}{\delta}\Big)^{2}>\frac{\nu+\delta}{\mu}=2$ and $\mathscr{R}_{0}^{\sigma}(\tfrac{b}{\delta},0)\approx 1.708$ so neither of the conditions for theorem 4.2 are met and we cannot guarantee stability in the stochastic case. In this case, the stochastic simulations often exhibit nontrivial outbreaks of the disease as seen in figure 4, where there is an initial outbreak of the disease among the noncompliant population, and later a significant outbreak among the compliant population that correlates with a random spike in noncompliance. Interestingly, in all realizations we observed, by time $T=50$, the stochastic simulation seems to match the deterministic one, which gives empirical evidence that the sufficient conditions for stability in theorem 4.2 may not be necessary (though it is not clear how they could be relaxed in our proof).
@@
For our third simulation, we consider $\beta=0.6,\xi=1$, $\mu=0.1$, $\nu=0$, and $\sigma_{\beta}=\sigma_{\mu}=0.4$. In this case, $\mathscr{R}_{0}^{\sigma}(0,\tfrac{b}{\delta})\approx 0.971$ and $0.16=\Big(\tfrac{\beta}{\delta}\Big)^{2}\sigma_{\mu}^{2}<\delta=0.2$, so the conditions of theorem 4.4 are met (as well as those of theorem 2.5(iii) in the worst case scenario of $\xi=1$ and $\nu=0$), so the disease free equilibrium $s=0,s^{*}=\tfrac{b}{\delta}$ is stable for both the deterministic and stochastic systems. This is seen in figure 5. If however, we increase $\sigma_{\mu}=2$ (leaving all other parameters as specified), we still have $\mathscr{R}_{0}^{\sigma}(0,\tfrac{b}{\delta})\approx 0.971$, but now $4=\Big(\tfrac{\beta}{\delta}\Big)^{2}\sigma_{\mu}^{2}>\delta=0.2$. Thus according to theorem 4.4, the infections should still exhibit exponential decay, but we can no longer guarantee stability for the DFE. This is seen in figure 6 where the infections do indeed die out, but there is some initial fighting between $S$ and $S^{*}$. As mentioned in remark 4.5, this would be unlikely to cause any concern for the policy-maker, since the important consideration is whether or not infections persist, not the particular concentration of compliance and noncompliance once infections have died out, and any additional compliant population is actually a boon since they will reduce the effective reproductive ratio.
@@
In our final simulation, we vary several of the parameters (including some that were fixed in figures 3-6). Specifically, we set the birth rate to $b=0.3$, the death rate to $\delta=1$, the infectivity rate of the disease to $\beta=0.5$, the recovery rate for the disease to $\gamma=0.25$, the infectivity rate for noncompliance to $\mu=8,$ the recovery rate from noncompliance to $\nu=1$, and the uncertainty in the infectivity rates to $\sigma_{\beta}=\sigma_{\mu}=0.125$. We take $\xi=0$ so that newly introduced members of the population are uniformly compliant. The reduction in infectivity is held at $\alpha=0.25$. With all these decisions, we consider the disease free state $s=\frac{\nu+\delta}{\mu}=\frac{1}{4}$ and $s^{*}=\tfrac{b}{\delta}-\frac{\nu+\delta}{\mu}=\frac{1}{20}$. In this case $\mathscr{R}_{0}(s,s^{*})\approx 0.0772$ so we are well within the local asymptotic stability regime provided by theorem 2.5(ii). Thus the solution of the deterministic system should tend toward the disease free equilibrium. While the stochastic system will not be at equilibrium at this pair $(s,s^{*})$, we do satisfy the hypothesis of theorem 4.7, and taking $C$ from (4.49), we have $C(\sigma_{\beta}^{2}+\sigma_{\mu}^{2})\approx 0.0171$. So while the solution of the stochastic system will not be steady, it should stay within these error bars of the deterministic solution. This is seen to occur in figure 7 where we plot the deterministic solution along with three realizations of the stochastic system for times $t\geq 10$. We see that the stochastic realizations stay well within the guaranteed error bars, which are displayed in gray.
@@ 6. Concluding remarks
In this manuscript we analyze ODE and stochastic ODE models for epidemiology incorporating human behavior in the form of noncompliance with NPIs. After analyzing DFEs for the deterministic model, we prove long time existence and positivity for the stochastic model and establish stability conditions for two DFEs by constructing suitable stochastic Lyapunov functions. Finally, we consider points which are DFEs of the deterministic model but not the stochastic model, and we quantify how much the stochastic model may stray from these points in the parameter regimes where they are stable for the deterministic model. We demonstrate our results with simulations.
From a epidemiological standpoint, our results indicate that, comparing a deterministic model to a stochastic one wherein infection rates (both for the disease and for the social contagion of noncompliant behavior) are modified with Gaussian white noise, the extinction criteria for the disease are strictly stronger in the stochastic case. Mathematically, this is seen in that sufficient conditions for disease extinction in theorems 4.2 and 4.4 are strictly stronger than their deterministic analogs. Likewise conditions which prevent the spread of noncompliance must be strengthened in the stochastic case. In some sense, this indicates that, while the white noise can either raise or lower infection rates, the policy-maker would need to plan for the worst.
We especially emphasize that theorems theorems 4.2 and 4.4 provide sufficient conditions for disease extinction. As such, they do not address disease persistence and may be able to relaxed. For example, in [22, 37] the authors derive thresholds for disease extinction or persistence wherein extinction becomes more likely in the presence of noise. However, the models in [22, 37] are somewhat simpler especially in that the dynamics of infected class have the form $dI=I\cdot(f(S,I)dt+g(S,I)dW)$, and this proportionality to $I$ is heavily exploited in the analysis. Because our model has competing nonlinearities and does not have this structure for the infected class dynamics, it is not straightforward to adapt the analysis [22, 37] to our case, and we instead have the stronger sufficient conditions for extinction as in [28, 51].
This work could be extended in many ways. One would be the identification of and stability conditions for endemic equilibrium points for the deterministic system and quantification of the behavior of stochastic system near these points. Another would be the incorporation of spatial effects resulting in a system of stochastic partial differential equation. There are some deterministic results in this direction in [9, 48] but adding stochasticity would lead to nontrivial complications in the analysis. Additionally, the inclusion of nonlocal effects wherein the disease and/or noncompliance can spread nonlocally could lead to interesting models requiring significantly modified analysis, and may be apt in the modeling of spread of attitudes, where social media, news coverage or a variety of other sources may be seen to influence behavior without direct contact. Another direction would be to explore the effects of introducing different types of stochastic perturbation, such as Stratonovich noise as in [7, 12, 19, 32], modified infections rates via Ornstein-Uhlenbeck processes as in [33, 38], or a mix of continuous- and discrete-time perturbations as in [24]. One last path forward would be to consider the governmental protocols as control variables and design a cost functional for the policy-maker so that this becomes a stochastic control problem.
@@ Acknowledgments
The authors would like to thank three anonymous reviewers for their insightful comments and remarks. WW was partially supported by a Junior Faculty Fellowship from the University of Oklahoma and a Simons Foundation grant (No. 0007730).